OpenStudy (anonymous):

a customer went to a garden shop and bought some potting soil for $12.50 and 6 shrubs. If the total bill was $71.00, how much did each shrub cost if they all cost the same? Write and solve an equation. Show your work.

OpenStudy (anonymous):

x=(71-12.50)/6 x=58.5/6 x=9.75 Each shrub cost $9.75